Olivia has just graduated from University and was hired by Manutech Inc. to help out with the companys financing decisions. The company has very high profit margins and generates large amounts of free cash flow. It currently has $500,000 in total assets and 10,000 shares outstanding. Because of generous investment tax credits and high rates of depreciation, Manutech does not pay any corporate tax. Manutech currently is all equity financed. The Chief Financial Officer ask Olivia to develop a financial strategy that could increase shareholder value. Olivia spends a week going over the companys financials, Olivia determines that there are three possible profit scenarios. Under the first scenario (recession), operating profits (EBIT) a S10,000 per year. Under the second scenario (expected), operating profits are $25,000 per year. Under the third scenario (expansion), operating profits are $40,000 per year re Olivia proposes the following plan, which will increase average earnings per share and return o equity. More specifically, she proposes that Manutech borrows money from a bank and use the proceeds to repurchase shares. As a result of this transaction, Manutech would have a debt-to- equity ratio of 2. The cost of borrowing from the bank is Please answer the following questions and document each step of your work (points will be deducted for not doing so). You must include a cover page with your submission.Question 3 (7.5 points) Calculate the break-even EBIT under which the ROE with 100% equity financing equals the ROE with a debt-to-equity ratio of 2. Plot the ROE versus EBIT for the different values of EBIT under the two financing arrangements (e.g. all equity versus debt financing).
Answer #1

1.

Break-even EBIT is the EBIT at which Return on Equity(ROE) is remain same, whether Firm is all equity financed or mix of equity and debt.

We have following information about Manutech Inc.

When all equity financed -

Total Assets = $ 500,000

No. of share = 10,000

When use debt-equity mix

Debt to equity = 2

Interest rate (i) = 3% p.a

Thus,

Debt = Total Assets*2/3 = 500,000*2/3

= $ 333,333

Equity = 500,000 - 333,333

= $ 166,667

Interest = $ 333,333*0.03

= $ 10,000

N etl ncome ROE = Equity

Where,

Net Income = EBIT - Interest - Tax

Manutech Inc. does not pay any taxes.

Thus,

Manutech's Net Income = EBIT - Interest.

In case of all equity financed, EBIT is equal to Net Income

Thus,

at Break-even EBIT

ROE of all equity financed = ROE of mix financed

EBIT EBIT- Interest Equity Equity

EBIT 500, 000 (EBIT - 10,000) 166,667

166, 667 * EBIT 500, 000 = (EBIT-10.000)

rac{EBIT}{3}=(EBIT-10,000)

EBIT=3EBIT-30,000

2EBIT=30,000

EBIT=rac{30,000}{2}

EBIT= 15,000

Thus, Break-even EBIT is $ 15,000.

2.

Plot of ROE vs. EBIT

Please refer to below sheet.

D17 All Equity financed 500,000 500,000 EBIT Total Assets 18% ROE In Recession Expected Expansion 10,000 25,000 40,000 2% 5% 8% 14% 10 Debt-Equity mix financed 10% 500,000 166,667 10,000 Total Assets 12 13 14 15 16 17 18 19 20 ul Interest EBIT ROE Recession Expected Expansion 10,000 25,000 40,000 18% 5,000 10,000 15,000 20,000 25,000 30,000 35,000 40,000 45,000 EBIT All Equity finance Debt-equity mix finance 23

Formula reference-

D17 A. All Equity financed Total Assets Equit In Recession Expected Expansion 18% EBIT ROE 10000 16% 10 Debt-Equity mix financed otal Assets l-C11/3 10000 12 13 14 15 16 Interest 6% EBIT ROE 4% Recession Expected Expansion C15-SC$13)/SC$12 C16-$C$13 (C17-$C$13)/$C$12 10000 12 40000 18 19 20 21 0 5,000 10,000 15,000 20,000 25,000 30,000 35,000 40,000 45,000 EBIT All Equity finance Debt-equity mix finance 23

Please note-

To Plot in excel, choose scatter with smooth lines and marker chart. Select EBIT for x-axis and ROE for y-axis.